OpenStudy (anonymous):

Help . The area of a rectangluar desk is given by the trinomial d^2 - 7d - 18. what are the possible dimensions of the desk? use factoring

jimthompson5910 (jim_thompson5910):

Find two numbers that multiply to -18 and add to -7

OpenStudy (anonymous):

9 and -2? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

close, -9 and 2

jimthompson5910 (jim_thompson5910):

-9*2 = -18 -9+2 = -7

jimthompson5910 (jim_thompson5910):

so d^2 - 7d - 18 factors to (d-9)(d+2)

jimthompson5910 (jim_thompson5910):

from the factorization, you can see the dimensions

OpenStudy (anonymous):

@jim_thompson5910 thank you :)

jimthompson5910 (jim_thompson5910):

np